OpenStudy (anonymous):

Solve. x2 + 10x = –21

OpenStudy (anonymous):

(x+3) (x+7)

OpenStudy (anonymous):

\[x^2+10x+21=0\] \[x^2+7x+3x+21=0\] \[x(x+7)+3(x+7)=0\] \[x+7=0,x+3=0\] \[x=-7,x=-3\]